# Personalized Negative Reservoir for Incremental Learning in Recommender Systems

### v1

@@ 4. Preliminaries
In this section, we provide the relevant background for our methodology. We succinctly review how a modern recommender system is trained using triplet loss and how knowledge distillation is applied in the incremental learning setting to alleviate forgetting.
4.1. Graph Based Recommendation Systems
@@
4.2. Knowledge Distillation for Inc. Learning
@@
Knowledge Distillation (KD) was originally designed to facilitate transferring the performance of a complex “teacher” model to a simpler “student” model (Hinton et al., 2015). This can be done by introducing an extra loss term to match the logits of the two models during training. In the setting of incremental learning for recommendation, the teacher model is the model trained on old data and the student model is trained on the most recent incremental block. The overall loss that we minimize takes the form:
(5) $\mathcal{L}_{BASE}=\mathcal{L}_{\textsc{BPR}}+\lambda_{\textsc{KD}}\mathcal{L}_{KD}({\mathcal{M}}_{T},{\mathcal{M}}_{S}),$
@@
where $\mathcal{L}_{\textsc{BPR}}$ is the BPR loss of the student model on the incremental data batch and $\mathcal{L}_{KD}$ represents the realization of the KD loss of teacher and student models, $\mathcal{M}_{T}$ and $\mathcal{M}_{S}$. The constant $\lambda_{\textsc{KD}}$ is the KD weight hyperparameter. Depending on the specific incremental learning technique, $\mathcal{L}_{KD}$ can take a different form and more components may be added to the overall loss function (Xu et al., 2020; Wang et al., 2020, 2021b).
5. Proposed Method: GraphSANE
@@
In this section we describe our proposed approach, the Graph Structure Aware NEgative (Graph-SANE) Reservoir for incremental learning of graph recommender systems. Our technique works by estimating the user interest shift with respect to item clusters between time blocks. It then uses the estimated user change of preferences to bias a negative sampler to provide high quality negative samples for the triplet loss. The following subsections describe how we construct, update and sample from our proposed personalized negative reservoir, and how the item clusters are obtained. We also present our overall training objective, which is end-to-end trainable. We note that our framework can be used by any graph neural network backbone and is compatible with existing incremental learning approaches. In this section we make no assumptions about the concrete algorithm realization besides the use of triplet loss.
5.1. Derivation
@@
Frequently, in a recommendation setting, we only have access to implicit feedback data. Concretely, this means that for each user $u$ we have access to the set of items that a user interacts with $\mathcal{I}_{u}^{+}=\{i:(u,i)\in\mathcal{E}\}$ but no explicit set of user dislikes $\mathcal{I}_{u}^{-}$. To optimize the model parameters $\Theta$, approaches such as the Bayesian Pairwise Loss (BPR) loss randomly samples a small number of items with which user $u$ has no observed interactions: $\mathcal{I}_{u}^{-}\subset\bar{\mathcal{I}}_{u}^{+}$, where ¯$\bar{\mathcal{I}}_{u}^{+}=\mathcal{I}\textbackslash\mathcal{I}_{u}^{+}$ represents the complementary set of $\mathcal{I}_{u}^{+}$. Thus, $\mathcal{I}_{u}^{-}$ is a randomly selected set of items that user $u$ does not interact with (Rendle et al., 2009). BPR samples each negative item once per user. However, in general, repeating some highly informative negatives multiple times (or weighting them more) can increase the speed of convergence. Furthermore, in the context of incremental learning the user interest shift can be used to drive a higher number of negatives from item categories that the user is losing interest in. The likelihood of a positive item $i$ being ranked above a negative item $j$ for user $u$ is:
(6) $p(\succ_{u}\mid\Theta)=\sigma(\hat{y}_{ui}-\hat{y}_{uj}).$
@@
Our proposed method proposes sampling negative items with replacement. Concretely, the proposed function $\textsc{SANE}_{\textsc{OPT}}$ with respect to which we optimize $\Theta$ may consider multiple copies $N_{u,j}$ of negative example $j$ for user $u$. This leads to $N_{u,j}\geq 0$ independent observations, each obeying the likelihood from (6):
@@
$\textsc{SANE}_{\textsc{OPT}}$ $\coloneqq\ln p(\Theta\mid\succ_{u})$
$\propto\ln\left(p(\succ_{u}\mid\Theta)p(\Theta)\right)$
@@
(7) $=\sum_{(u,i)\in\mathcal{E},j\in\mathcal{I}_{u}^{-}}N_{u,j}\ln\sigma(\hat{y}_{ui}-\hat{y}_{uj})-\lambda_{\Theta}\|\Theta\|^{2}.$
In practice we minimize the negative of (5.1) using backpropagation:
(8) $\mathcal{L}_{\textsc{SANE}}=-\textsc{SANE}_{\textsc{OPT}}.$
Our loss is now defined. In the next section we demonstrate a concrete procedure to obtain $(N_{u,1},\dots,N_{u,|\mathcal{I}|})$ for each user.
5.2. Negative Reservoir for Inc. Learning
@@
Determining user interest shift: We propose tracking user interests by measuring the number of interactions of each user with item categories at every time step. For example, in the toy example depicted in Fig. 1, we see that there are $K=3$ item categories. By counting the proportion of items that each user interacts with (in Fig. 1 (b)) we obtain histograms of user-item category interactions $\mathbf{H}_{u,t}\in\mathbb{R}^{K}$. These histograms are then normalized and projected to the simplex $\Delta_{K-1}$ (in Fig. 1 (c)). By tracking the trajectory of each user on the simplex we can surmise the user’s interest shift:
(9) $\text{shift}(u,t-1,t)=\frac{\mathbf{H}_{u,t}}{|\mathbf{H}_{u,t}|}-\frac{\mathbf{H}_{u,t-1}}{|\mathbf{H}_{u,t-1}|},$
where ||$\cdot$|| ⋅ | denotes the L1 norm. Note that our method is robust to the case where item categories are not part of the item feature set. In this case we cluster the items and interpret the item clusters as induced pseudo-categories.
@@
Negative items: Consider the top ranked $Q$ negative items for user $u$ (by sorting row $u$ of $\mathbf{R}$) at time $t$. That is the set of size $Q$ that contains the highest ranked items the our model produces for user $u$ that he/she does not interact with. For the top $Q$ negative items per user we note the item category (or item cluster membership if categories are not available in the data) of each item. We denote $K$ as the number of item categories/clusters. This yields a count of top negative item interactions per item category for each user $\mathbf{c}_{u,t}=(\mathbf{c}_{u,t,1},\dots,\mathbf{c}_{u,t,K})$, such that $\sum_{i=1}^{K}\mathbf{c}_{u,t,i}=Q$. For example, in Fig. 1 (d), the vector would contain the counts of negative pet, car and coffee items associated with a particular user at a specific time step. We interpret the observed top negative user item-category interactions $\mathbf{c}_{u,t}\in\mathbb{N}^{K}=(c_{u,t,1},\dots,c_{u,t,K})$ as $Q$ samples of a multinomial distribution with unknown parameters $\bm{\theta}_{u,t}\in\mathbb{R}^{K}=(\theta_{u,t,1},\dots,\theta_{u,t,K})$.
@@
Our hypothesis: A good sampling distribution for negatives should prioritize sampling negative interactions that simultaneously: (i) correspond to item categories for which the user exhibits reduced interest at the present time when compared to his/her historical preferences, and (ii) are ranked highly (hard negatives).
@@
Hierarchical Model: We model the prior distribution of $\bm{\theta}_{u,t}$ as a Dirichlet $\bm{\theta}_{u,t}\sim\textbf{Dir}(\bm{\alpha}_{u,t})$. Our framework obtains a posterior distribution over $\bm{\theta}_{u,t}$ by fusing information from our prior model with the observed distribution of negative items across categories. Looking at the positive user item interactions, the parameters of the prior $\bm{\alpha}_{u,t}$ are derived from the histogram of user item category interaction counts $\mathbf{H}_{u,t}$ and $\mathbf{H}_{u,t-1}$. $\mathbf{H}_{u,t}\in\mathbb{N}^{K}$ is a histogram that summarizes the number of items that user $u$ interacted with items from each of the $K$ categories, i.e., $\mathbf{H}_{u,t}[k]=n$ denotes that at time $t$ user $u$ interacted with $n$ items from category $k$. Recall that $Q$ is the number of negative items in each user’s negative reservoir. We then define the following hierarchical model (summarized in Fig. 2):
@@
(10) LEVEL 1: $\bm{\alpha}_{u,t}=\lambda Q\operatorname{\textsc{softmax}}\bigg{(}\frac{\mathbf{H}_{u,t}}{|\mathbf{H}_{u,t}|}-\frac{\mathbf{H}_{u,t-1}}{|\mathbf{H}_{u,t-1}|}\bigg{)}\,,$
@@
where ||$\cdot$|| ⋅ | denotes the L1 norm and $\bm{\alpha}_{u,t}\in\mathbb{R}^{K}$. Our choice for prior in (10) reflects the fact that if for the $k$-th index of $[\bm{\alpha}_{u,t}]_{k}=\left[\frac{\mathbf{H}_{u,t}}{|\mathbf{H}_{u,t}|}-\frac{\mathbf{H}_{u,t-1}}{|\mathbf{H}_{u,t-1}|}\right]_{k}<0$ then a user is interacting with fewer items belonging to category $k$, and hence may be losing interest in the item category with index $k$, so we want to sample more negative items from it. Conversely, if the difference is positive, this indicates an increase in user interest in items from the category $k$ so the softmax function will decrease the probability of sampling a negative item from this item category. The next levels are a standard Multinomial-Dirichlet conjugate pair:
@@
(11) LEVEL 3: $f_{u,t}(\mathbf{c}_{u,t}\mid\bm{\theta}_{u,t})=\frac{N!}{\prod_{i=1}^{K}\mathbf{c}_{u,t,i}!}\prod_{i=1}^{K}\theta_{u,t,i}^{\mathbf{c}_{u,t,i}}\,.$
@@
$p(\bm{\theta}_{u,t}\mid\mathbf{c}_{u,t},\bm{\alpha}_{u,t})\propto p(\bm{\theta}_{u,t})f_{u,t}(\mathbf{c}_{u,t}\mid\mathbf{\theta}_{u,t})$
@@
$=\frac{\Gamma(\alpha_{0})}{\Gamma(\alpha_{1})\dots\Gamma(\alpha_{K})}\prod_{k=1}^{K}\theta_{k}^{a_{k}-1}\frac{N!}{\prod_{i=1}^{K}c_{i}!}\prod_{i=1}^{K}\theta_{i}^{c_{i}}=\text{{Dir}}(\bm{\alpha}_{u,t}+\mathbf{c}_{u,t})$
(12) $=\text{{Dir}}\bigg{(}\lambda Q\operatorname{\textsc{softmax}}\bigg{(}\frac{\mathbf{H}_{u,t}}{|\mathbf{H}_{u,t}|}-\frac{\mathbf{H}_{u,t-1}}{|\mathbf{H}_{u,t-1}|}\bigg{)}+\mathbf{c}_{u,t}\bigg{)}$
Figure 2. The hierarchical model in (10), (5.2). From time $t-1$ to $t$, the user loses interest in the i-th category, this increases $\bm{\alpha}_{i}$. This gives $\bm{\theta}_{i}$ a high prior value, thus we sample negatives with higher probability from that item category in (14).
We can then estimate $\bm{\theta}$ as the mean of the posterior:
@@
(13) $\bm{\hat{\theta}}_{i}=\mathbb{E}[\mathbf{\theta}_{i}\mid\mathbf{c}_{u,t,i}]$ $=\frac{\lambda Q\operatorname{\textsc{softmax}}\bigg{(}\frac{\mathbf{H}_{u,t,i}}{|\mathbf{H}_{u,t,i}|}-\frac{\mathbf{H}_{u,t-1,i}}{|\mathbf{H}_{u,t-1,i}|}\bigg{)}+\mathbf{c}_{u,t,i}}{\sum_{i}\lambda Q\operatorname{\textsc{softmax}}\bigg{(}\frac{\mathbf{H}_{u,t,i}}{|\mathbf{H}_{u,t,i}|}-\frac{\mathbf{H}_{u,t-1,i}}{|\mathbf{H}_{u,t-1,i}|}\bigg{)}+\mathbf{c}_{u,t,i}}$
@@
Sampler for Reservoir: Denote $\bm{\hat{\theta}}_{u,t,C}$ as the $C$-th entry of $\bm{\hat{\theta}}_{u,t}$ ($C\in\{1,\dots,K\}$ is the item category that $i$ belongs to). Now, to sample negative items from the reservoir, we define the following sampler that draws negative item $n_{i}\in\mathcal{I}$ according to:
(14) $p(n_{i}=i\mid i\in C,u,\bm{\hat{\theta}}_{u,t})\vcentcolon=\frac{\mathbb{I}_{[i\in\mathcal{Q}]}(i)\bm{\hat{\theta}}_{u,t,C}}{\sum_{i}\mathbb{I}_{[i\in\mathcal{Q}]}(i)\bm{\hat{\theta}}_{u,t,C}},$

### v2

@@ 4 Preliminaries
In this section, we provide the relevant background for our methodology. We succinctly review how a modern recommender system is trained using triplet loss and how knowledge distillation is applied in the incremental learning setting to alleviate forgetting. This section also serves to introduce notation.
4.1 Graph Based Recommendation Systems
@@
4.2 Knowledge Distillation for Incremental Learning
@@
Knowledge Distillation (KD) was originally designed to facilitate transferring the performance of a complex “teacher” model to a simpler “student” model (Hinton et al., 2015). In the setting of incremental learning for recommendation, the teacher model is the model trained on old data and the student model is trained on the most recent incremental block. The overall loss that we minimize takes the form:
$\mathcal{L}_{BASE}=\mathcal{L}_{\textsc{TRIPLET}}+\lambda_{\textsc{KD}}\mathcal{L}_{KD}({\mathcal{M}}_{T},{\mathcal{M}}_{S}),$ (4)
@@
where $\mathcal{L}_{\textsc{TRIPLET}}$ is a standard triplet loss for recommendation such as the BPR loss (Rendle et al., 2009) (defined in the Appendix B) of the student model on the incremental data batch and $\mathcal{L}_{KD}$ represents the realization of the KD loss of teacher and student models, $\mathcal{M}_{T}$ and $\mathcal{M}_{S}$. The constant $\lambda_{\textsc{KD}}$ is the KD weight hyperparameter. Depending on the specific incremental learning technique, $\mathcal{L}_{KD}$ can take a different form and more components may be added to the overall loss function (Xu et al., 2020; Wang et al., 2020; 2021b). For concrete examples of realizations of KD losses in our setting see the Appendix B.
5 Proposed Method: GraphSANE
@@
In this section we describe our proposed approach, the Graph Structure Aware NEgative (Graph-SANE) Reservoir for incremental learning of graph recommender systems. Our technique works by estimating the user interest shift with respect to item clusters between time blocks. It then uses the estimated user change of preferences to bias a negative sampler to provide high quality negative samples for the triplet loss. The following subsections describe how we construct, update and sample from our proposed personalized negative reservoir, and how the item clusters are obtained. We also present our overall training objective, which is end-to-end trainable. We note that our framework can be used by any graph neural network backbone that uses triplet loss and is compatible with existing incremental learning approaches.
On a high level, our method consists of three components. First, we introduce the negative reservoir in Section 5.1 where we define loss term $\mathcal{L}_{\textsc{SANE}}$. We detail how to update the contents of this reservoir for each user in Section 5.2. The user interest shift is computed based on clustering the items the user interacts with. One such end-to-end trainable clustering technique (with loss component $\mathcal{L}_{\textsc{KL}}$) from the literature is reviewed in Section 5.3. Finally, Section 5.4 where the incremental learning negative reservoir loss $\mathcal{L}_{\textsc{SANE}}$ and differentiable clustering loss $\mathcal{L}_{\textsc{KL}}$ is combined with the base model loss $\mathcal{L}_{\textsc{BASE}}$ to produce the overall objective that is optimized via backpropagation.
5.1 Derivation
@@
To optimize the model parameters $\Theta$, triplet loss approaches randomly sample a small number of items with which user $u$ has no observed interactions: $\mathcal{I}_{u}^{-}\subset\bar{\mathcal{I}}_{u}^{+}$, where ¯$\bar{\mathcal{I}}_{u}^{+}=\mathcal{I}\textbackslash\mathcal{I}_{u}^{+}$ represents the complementary set of $\mathcal{I}_{u}^{+}$. Thus, $\mathcal{I}_{u}^{-}$ is a randomly selected set of items that user $u$ does not interact with (Rendle et al., 2009). Existing approaches sample each negative item once per user. However, in general, repeating some highly informative negatives multiple times (or weighting them more) can increase the speed of convergence. Furthermore, in the context of incremental learning the user interest shift can be used to drive a higher number of negatives from item categories that the user is losing interest in. The likelihood of a positive item $i$ being ranked above a negative item $j$ for user $u$ is:
$p(\succ_{u}\mid\Theta)=\sigma(\hat{y}_{ui}-\hat{y}_{uj}).$ (5)
@@
Our proposed method proposes sampling negative items with replacement. Concretely, the proposed function $\textsc{SANE}_{\textsc{OPT}}$ with respect to which we optimize $\Theta$ may consider multiple copies $N_{u,j}$ of negative example $j$ for user $u$. This leads to $N_{u,j}\geq 0$ independent observations, each obeying the likelihood from equation 5:
@@
$\textsc{SANE}_{\textsc{OPT}}$ $\coloneqq\ln p(\Theta\mid\succ_{u})\propto\ln\left(p(\succ_{u}\mid\Theta)p(\Theta)\right)$
@@
$=\sum_{(u,i)\in\mathcal{E},j\in\mathcal{I}_{u}^{-}}N_{u,j}\ln\sigma(\hat{y}_{ui}-\hat{y}_{uj})-\lambda_{\Theta}\|\Theta\|^{2},$ (6)
where $p(\Theta)$ is an isotropic normal distribution. In practice we minimize the negative of equation 5.1 using backpropagation: $\mathcal{L}_{\textsc{SANE}}=-\textsc{SANE}_{\textsc{OPT}}.$ Our loss is now defined. In the next section we demonstrate a concrete procedure to obtain $(N_{u,1},\dots,N_{u,|\mathcal{I}|})$.
5.2 Negative Reservoir for Incremental Learning
Figure 1: Toy example with daily model updates and 3 item categories: pets, car parts and coffee items. The figure shows a user’s interactions with these categories over a week. (a) Cumulative user interest over time: A larger fraction of user $u$’s interactions comes from pet items at the start of the week, but his interests change over time and he interacts with more car part items by the end of the week. (b) The change of preference is reflected on the daily histograms of user-item-category interactions $\mathbf{H}_{u,t}$. (c) The histograms are normalized $\widehat{\mathbf{H}}_{u,t}$ and projected on the simplex $\Delta_{2}$. There we see the interest shift on the simplex representing the user’s interest shift from $t=2,$ (Tuesday) to $t=3$, (Wednesday) and on the right the user’s interest shift from $t=3,$ (Wednesday) to $t=4$, (Thursday). We see that the user exhibits a large change in interests on the second simplex (moves far away from current neighborhood in simplex denoted by a red dashed circle). (d) We propose to sample more negatives from pet category when fine tuning the model on Thursday to allow the model to quickly adjust to new user interests (in the figure the right bucket has many more pet items).
@@ 5.2 Negative Reservoir for Incremental Learning
Determining user interest shift: We propose tracking user interests by measuring the number of interactions of each user with item categories at every time step. For example, in the toy example depicted in Fig. 1, we see that there are $K=3$ item categories. By counting the proportion of items that each user interacts with (in Fig. 1 (b)) we obtain histograms of user-item category interactions $\mathbf{H}_{u,t}\in\mathbb{R}^{K}$. These histograms are then normalized and projected to the simplex $\Delta_{K-1}$ (in Fig. 1 (c)). By tracking the trajectory of each user on the simplex we can surmise the user’s interest shift: $\text{shift}(u,t-1,t)=\frac{\mathbf{H}_{u,t}}{|\mathbf{H}_{u,t}|}-\frac{\mathbf{H}_{u,t-1}}{|\mathbf{H}_{u,t-1}|},$ where ||$\cdot$|| ⋅ | denotes the L1 norm. Note that our method is robust to the case where item categories are not given. In this case we cluster the items and interpret the item clusters as induced pseudo-categories.
Figure 2: The hierarchical model in equation 7 & 8. From time $t-1$ to $t$, a user loses interest in the i-th category, this increases $\bm{\alpha}_{i}$. Thus $\bm{\theta}_{i}$ has a high prior value so we sample negatives with higher probability from that item category in equation 11.
@@
Negative items: Consider the top ranked $Q$ negative items for user $u$ (by sorting row $u$ of $\mathbf{R}$) at time $t$. This is the set of size $Q$ that contains the items that the model ranks the highest for user $u$, but which the user does not interact with. For the top $Q$ negative items per user we note the item category (or item cluster membership if categories are not available in the data) of each item. We denote $K$ as the number of item categories/clusters. This yields a count of top negative item interactions per item category for each user $\mathbf{c}_{u,t}=(\mathbf{c}_{u,t,1},\dots,\mathbf{c}_{u,t,K})$, such that $\sum_{i=1}^{K}\mathbf{c}_{u,t,i}=Q$. For example, in Fig. 1 (d), this vector would contain the counts of negative pet, car and coffee items associated with a particular user at a specific time step. We interpret the observed top negative user item-category interactions $\mathbf{c}_{u,t}\in\mathbb{N}^{K}=(c_{u,t,1},\dots,c_{u,t,K})$ as $Q$ samples of a multinomial distribution with unknown parameters $\bm{\theta}_{u,t}\in\mathbb{R}^{K}=(\theta_{u,t,1},\dots,\theta_{u,t,K})$, where $\theta_{u,t,K}$ represents the probability of sampling a negative item from category $K$ for user $u$ at time $t$.
@@
Our hypothesis: A good sampling distribution for negatives should prioritize sampling negative interactions that simultaneously: (i) correspond to item categories for which the user exhibits reduced interest at the present time when compared to his/her historical preferences; and (ii) are ranked highly (hard negatives)
@@
Hierarchical Model: We model the prior distribution of $\bm{\theta}_{u,t}$ as a Dirichlet $\bm{\theta}_{u,t}\sim\textbf{Dir}(\bm{\alpha}_{u,t})$. Our framework obtains a posterior distribution over $\bm{\theta}_{u,t}$ by fusing information from our prior model with the observed distribution of negative items across categories. The parameters of the prior $\bm{\alpha}_{u,t}$ are derived from the histogram of user item-category interaction counts $\mathbf{H}_{u,t}$ and $\mathbf{H}_{u,t-1}$. $\mathbf{H}_{u,t}\in\mathbb{N}^{K}$ is a histogram that summarizes the number of items that user $u$ interacted with from each of the $K$ categories, i.e., $\mathbf{H}_{u,t}[k]=n$ denotes that at time $t$ user $u$ interacted with $n$ items from category $k$. Recall that $Q$ is the number of negative items in each user’s negative reservoir. We then define the following hierarchical model (shown in Fig. 2):
@@
LEVEL 1: $\bm{\alpha}_{u,t}=\,\lambda\,Q\,\mathrm{softmax}\bigg{(}-\left[\frac{\mathbf{H}_{u,t}}{|\mathbf{H}_{u,t}|}-\frac{\mathbf{H}_{u,t-1}}{|\mathbf{H}_{u,t-1}|}\right]\bigg{)}\,,$ (7)
@@
where ||$\cdot$|| ⋅ | denotes the L1 norm and $\bm{\alpha}_{u,t}\in\mathbb{R}^{K}$. Our choice for prior in equation 7 reflects the fact that if $\left[\frac{\mathbf{H}_{u,t}}{|\mathbf{H}_{u,t}|}-\frac{\mathbf{H}_{u,t-1}}{|\mathbf{H}_{u,t-1}|}\right]_{k}<0$, then a user is interacting with fewer items belonging to category $k$, and hence may be losing interest in the item category with index $k$, so we want to sample more negative items from it. Conversely, if the difference is positive, this indicates an increase in user interest in items from the category $k$ so the softmax function will decrease the probability of sampling a negative item from this category. The next levels are a Multinomial-Dirichlet conjugate pair:
@@
LEVEL 3: $f_{u,t}(\mathbf{c}_{u,t}\mid\bm{\theta}_{u,t})=\frac{N!}{\prod_{i=1}^{K}\mathbf{c}_{u,t,i}!}\prod_{i=1}^{K}\theta_{u,t,i}^{\mathbf{c}_{u,t,i}}\,.$ (8)
@@
$p(\bm{\theta}_{u,t}\mid\mathbf{c}_{u,t},\bm{\alpha}_{u,t})\propto p(\bm{\theta}_{u,t})f_{u,t}(\mathbf{c}_{u,t}\mid\mathbf{\theta}_{u,t})$ $=\frac{\Gamma(\alpha_{0})}{\Gamma(\alpha_{1})\dots\Gamma(\alpha_{K})}\prod_{k=1}^{K}\theta_{k}^{a_{k}-1}\frac{N!}{\prod_{i=1}^{K}c_{i}!}\prod_{i=1}^{K}\theta_{i}^{c_{i}}=\text{{Dir}}(\bm{\alpha}_{u,t}+\mathbf{c}_{u,t})$
$=\text{{Dir}}\bigg{(}\lambda Q\mathrm{softmax}\bigg{(}\frac{\mathbf{H}_{u,t}}{|\mathbf{H}_{u,t}|}-\frac{\mathbf{H}_{u,t-1}}{|\mathbf{H}_{u,t-1}|}\bigg{)}+\mathbf{c}_{u,t}\bigg{)}$ (9)
We can then estimate $\bm{\theta}_{u,t}$ as the mean of the posterior:
@@
$\bm{\hat{\theta}}_{u,t,i}=\mathbb{E}[\mathbf{\theta}_{u,t,i}\mid\mathbf{c}_{u,t,i}]$ $=\frac{\lambda Q\mathrm{softmax}\bigg{(}\frac{\mathbf{H}_{u,t,i}}{|\mathbf{H}_{u,t,i}|}-\frac{\mathbf{H}_{u,t-1,i}}{|\mathbf{H}_{u,t-1,i}|}\bigg{)}+\mathbf{c}_{u,t,i}}{\sum_{i}\lambda Q\mathrm{softmax}\bigg{(}\frac{\mathbf{H}_{u,t,i}}{|\mathbf{H}_{u,t,i}|}-\frac{\mathbf{H}_{u,t-1,i}}{|\mathbf{H}_{u,t-1,i}|}\bigg{)}+\mathbf{c}_{u,t,i}}$ (10)
@@
Sampler for Reservoir: Denote $\bm{\hat{\theta}}_{u,t,C}$ as the $C$-th entry of $\bm{\hat{\theta}}_{u,t}$ ($C\in\{1,\dots,K\}$ is the item category that $i$ belongs to). To sample negative items from the reservoir, we define the sampler for negative item $n_{i}\in\mathcal{I}$:
$p(n_{j}=j,\bm{\hat{\theta}}_{u,t})\vcentcolon=\frac{\mathbb{I}_{[j\in\mathcal{Q}]}(j)\bm{\hat{\theta}}_{u,t,g(j)}}{\sum_{i}\mathbb{I}_{[i\in\mathcal{Q}]}(i)\bm{\hat{\theta}}_{u,t,g(i)}},$ (11)